Problem 3

At an annual percentage rate (APR) of 16%, compounded quarterly, what is the present value of an income stream specified as follows? It pays $60 next quarter and grows at 2% per quarter until the end of the third year, and from that point on it grows at 1% per quarter indefinitely
Answer #1

- 100 101 Formatting Styles 2 Clear Format Painter Clipboard A B as Table Styles Font D Alignment H C Filter Editing R E F G

A C D E F G I J K L M N O P Q 2 Quarterly -16%/4 Year 1 Year 2 Year 3 Year 01 2 3 8 10 11 12 Cash flow 60 =E6*1.02 =F6*1.02 =